OpenStudy (anonymous):

Write the equation of the line perpendicular to the line x - 3y = 9 and passing through the point (3,5). Write the equation of the line parallel to the line 10x - 5y = 8 and passing through the point (2,4). Write the equation of the line parallel to the line 4x + y = -1 and passing through the point (5,0). Write the equation of the line perpendicular to the line x - 5y = -10 and passing through the point (2,5). Write the equation of the line parallel to the line 4x + 2y = 5 and passing through the point (-3,5).

OpenStudy (anonymous):

A) y = -4x + 20 B) y = -2x - 1 C) y = 2x D) y = -5x + 15 E) y = -3x + 14

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

@myininaya @TuringTest

OpenStudy (anonymous):

can anyone help mee!!

OpenStudy (anonymous):

i can help you

OpenStudy (anonymous):

@tkhunnу @tkhunny

OpenStudy (anonymous):

forget it i guessed!

OpenStudy (tkhunny):

How do you build the equation of a line? What information do you need? How do you get the necessary information from the given information?

OpenStudy (anonymous):

idk

OpenStudy (tkhunny):

That is a gravely insufficient response unless your goal is to fail. You've NEVER heard of the "Point-Slope" form of a line? You've NEVER heard of the "2-Point" form of a line? You've NEVER heard of the "Slope-Intercept" form of a line? Is this what you mean by "I Don't Know"? Please confirm.

OpenStudy (anonymous):

yea tk he knows it all he sits next to meh

OpenStudy (tkhunny):

Well, there you go. In each case, you have a point and a hint to find the slope. Go!